OpenStudy (anonymous):

how do i find the volume of a cone where the height is 2 square root 17 and the radius is 3 square root 2

OpenStudy (anonymous):

pleaase help!!!!!!

jhonyy9 (jhonyy9):

do you know formula for volum of a cone ?

jhonyy9 (jhonyy9):

V= area of base time height divide by 3 V = pi *r^2 *h /3 V = pi*18*2sqrt17 / 3 = 6pi*2sqrt17 = 12pisqrt17 V = 12pi*sqrt17